OpenStudy (anonymous):

find the equation of the line with x-intercept 4 and y-intercept 8

OpenStudy (anonymous):

first find slope, y2 - y1 / x2 - x1 so for you, you can get the points (0,8) and (4,0) plug that in and you get 0 - 8 / 4 - 0 giving you a slope of -2 (-8/4) if your line is going the other way (from (4,0) to (0,8) then it would be -1/2) to find the last value all you have to do is plug in one of your x values and find the needed adjustment to make it equal the corresponding y value. so take y = -2x and plug in 0 for x you get y = -2(0) --> y = 0 to make y = 8 then you must add 8 finally you need to double check, so do y = -2x + 8 but plug in the other set of points 0 = -2(4) +8 0 = -8 + 8 0 = 0 so we know y = -2x + 8 is correct

OpenStudy (anonymous):

thank u